The coordinates of a particle moving in xy plane at any time t are 2t,t^2 . Find trajectory of the particle. 

Grade:12

2 Answers

Aditya Gupta
4 years ago
hello sulipsa this is an easy ques and can be done as follows:
we are given that x= 2t and y= t^2
so to find the trajectory of the particle, we need to find a relation b/w x and y coordinates.
so we need to eliminate t.
we see that y= t^2 or 4y= (2t)^2
but 2t= x, so
4y= x^2, which is an upward facing parabola.
kindly approve :)
Khimraj
4 years ago
 

Given co-ordinate of the particle as ( 2t,t^2)

means at ny time  t....x = 2t   and y = t^2

 

i.e t = x/2     and so  y = ( x/2)^2   or   4y = x^2

Hence the trajectory of the particle is a parabola with equation  x^2 = 4y .....